OpenStudy (anonymous):

What are the discontinuities of the function f(x) = the quantity of x squared plus 6 x plus 9, all over 3 x plus 15.?

OpenStudy (anonymous):

medal to whoever gives me the answer. and it be right

OpenStudy (freckles):

just find for what x the bottom is zero

OpenStudy (anonymous):

@mathmath333 @princeharryyy @satellite73 @DullJackel09

OpenStudy (freckles):

that is all you have to this since you have a polynomial over a polynomial is just take that bottom polynomial and set it equal to 0 and solve for x

OpenStudy (freckles):

solve: 3x+15=0

OpenStudy (anonymous):

so 5?

OpenStudy (anonymous):

-5*

OpenStudy (freckles):

yes 3(-5)+15=-15+15=0 so yes x=-5 is when the bottom is 0